Thermal Rising, Inc., makes paragliders for sale through specialty sporting goods stores. The company has a standard paraglider model, but also makes custom-designed paragliders. Management has designed an activity-based costing system with the following activity cost pools and activity rates:

Activity Cost Pool Activity Rate
Supporting direct labor $ 22 per direct labor-hour
Order processing $ 200 per order
Custom design processing $ 260 per custom design
Customer service $ 424 per customer

Management would like an analysis of the profitability of a particular customer, Big Sky Outfitters, which has ordered the following products over the last 12 months:

Standard
Model
Custom
Design
Number of gliders 11 2
Number of orders 2 2
Number of custom designs 0 2
Direct labor-hours per glider 29.50 33.00
Selling price per glider $ 1,650 $ 2,400
Direct materials cost per glider $ 446 $ 566

The company’s direct labor rate is $16 per hour.

Required:

Using the company’s activity-based costing system, compute the customer margin of Big Sky Outfitters. (Round your intermediate calculations and final answer to the nearest whole dollar amount. Loss amounts should be entered with a minus sign.)
Answer #1
Revenue 22950 =(1650*11)+(2400*2)
Costs:
Direct materials 6038 =(446*11)+(566*2)
Direct labor 6248 =(29.5*16*11)+(33*16*2)
Supporting Direct labor 8591 =(29.5*22*11)+(33*22*2)
Order processing 800 =(2*200)+(2*200)
Custom design processing 520 =(2*260)
Customer service 424
Total costs 22621
Customer margin 329